# Deep Networks Always Grok and Here is Why

### v1

@@ 4 What Affects the Progress Measure?
Parameterization In Figures 12, 19 and 21, we see that increasing the number of parameters either by increasing depth, or by increasing width of the network in our MNIST-MLP experiments, hastens region migration, therefore makes grokking happen earlier.
@@ 5 Conclusions and Limitations
We persued a thorough empirical study of grokking, both on the test dataset and adversarial examples generated using the test dataset. We obtained new observations hinting that grokking is a common phenomenon in deep learning that is not restricted to particular tasks or DNN initialization. Upon this discovery, we delved into DNNs geometry to isolate the root cause of both delayed generalization and robustness which we attributed to the DNN’s linear region migration that occurs in the latest phase of training. Again, the observation of such migration of the DNN partition is a new discovery of its own right. We hope that our analysis has provided novel insights into DNNs training dynamics from which grokking naturally emerges. While we empirically study the local complexity dynamics, a theoretical justification behind the double descent behavior is lacking. At a high level, it is clear that the classification function being learned has its curvature concentrated at the decision boundary and approximation theory would normally dictate a free-form spline to therefore concentrate its partition regions around the decision boundary to minimize approximation error. However, it is not clear why that migration occurs so late in the training process, and we hope to study that in future research. We also see empirical evidence of region migration while using Adam as the optimizer. The training dynamics of stochastic gradient descent, as well as sharpness aware minimization (Andriushchenko & Flammarion, 2022) can also be studied using our framework. There can be possible connections between region migration and neural collapse (Papyan et al., 2020) which are not explored in this paper. The spline viewpoint of deep neural networks may provide strong geometric insights to assist in mechanistic understanding in future works as well.
@@ Appendix D Extra Figures
Layer 1Refer to caption
Layer 2Refer to caption
Layer 3Refer to caption
Layer 4Refer to caption
Layer 5Refer to caption
@@
Figure 24: Dataset size does not affect the onset of region migration. Local complexity dynamics training an MLP on MNIST with weight decay

### v2

@@ 4 What Affects the Progress Measure?
Parameterization. In Figures 14, 27 and 29, we see that increasing the number of parameters either by increasing depth, or by increasing width of the network in our MNIST-MLP experiments, hastens region migration, therefore makes grokking happen earlier.
@@ 5 Conclusions and Limitations
We have pursued a thorough empirical study of grokking, both on the test dataset and adversarial examples generated using the test dataset. We obtained new observations hinting that grokking is a common phenomenon in deep learning that is not restricted to particular tasks or DNN initialization. Upon this discovery, we delved into DNNs geometry to isolate the root cause of both delayed generalization and robustness which we attributed to the DNN’s linear region migration that occurs in the latest phase of training. Again, the observation of such migration of the DNN partition is a new discovery of its own right. We hope that our analysis has provided novel insights into DNNs training dynamics from which grokking naturally emerges. While we empirically study the local complexity dynamics, a theoretical justification behind the double descent behavior is lacking. At a high level, it is clear that the classification function being learned has its curvature concentrated at the decision boundary and approximation theory would normally dictate a free-form spline to therefore concentrate its partition regions around the decision boundary to minimize approximation error. However, it is not clear why that migration occurs so late in the training process, and we hope to study that in future research. We also see empirical evidence of region migration while using Adam as the optimizer. The training dynamics of stochastic gradient descent, as well as sharpness aware minimization (Andriushchenko & Flammarion, 2022) can also be studied using our framework. There can be possible connections between region migration and neural collapse (Papyan et al., 2020) which are not explored in this paper. The spline viewpoint of deep neural networks may provide strong geometric insights to assist in mechanistic understanding in future works as well.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. One takeaway of this work is that training Deep Neural Networks longer may lead to increased robustness. Training networks especially foundation models for longer may have potential societal consequences in terms of carbon emissions.
@@ Appendix A Empirical analysis of our proposed method
Optimization Steps
Figure 16: Training accuracy and robust accuracy for the networks trained on randomly labeled MNIST samples presented in Figure 13.
Figure 17: Increasing the batch-size expedites grokking. This indicates that reduced SGD noise allows region migration to occur earlier in training.
Figure 18: Grokking stronger adversarial attacks. We see that during delayed generalization, the robustness to Auto-Attack (Croce & Hein, 2020) also increases. This shows the universality of delayed robustness.
Figure 19: Token embedding space LC on a 2D subspace intersecting three random training points. We visualize the LC layerwise for the GeLU activated MLP layers inside each of the 12 blocks of the LLM for which we present training dynamics in Figure 9. The LC is computed after 197 optimization steps, during the peak ascent. We see that LC on this subspace is very high especially close to the data points. LC values are clamped to a maximum of 150.
Figure 20: Token embedding space LC on a 2D subspace intersecting three random training points. We visualize the LC layerwise for the GeLU activated MLP layers inside each of the 12 blocks of the LLM for which we present training dynamics in Figure 9. The LC is computed after 372759 optimization steps, therefore during the second LC descent. We see that LC on this subspace is concentrated away from the training points, especially for the deeper layers. This indicates that region migration occurs in LLMs as well, leading to delayed robustness. LC values are clamped to a maximum of 150.
@@ Appendix D Extra Figures
Layer 1
Layer 2
Layer 3
Layer 4
Layer 5
@@
Input Partition
LC, $r=0.05$
LC, $r=0.1$
LC, $r=0.5$
LC, $r=1$
Figure 25: Input space partition computed analytically via SplineCam (Humayun et al., 2023b) for the 2D toy setting presented in Figure 3 (left). Regions are colored by white and knots are colored by red. The partition is computed for the input space domain $[-10,10]^{2}$, induced by an MLP of depth $5$ and width $30$. We take a meshgrid of $300\times 300$ points over the input domain, and measure the local complexity at each point with radius, $r\in\{0.05,0.1,0.5,1\}$ (rest). We see that our proposed method can locate the non-linearities for small $r$. As $r$ is increased our method provides a coarser estimate of the local density of non-linearities, i.e., number of non-linearities intersecting the a fixed volume defined by the local neighborhood.
@@
Figure 32: Increasing training data size expedites region migration. Local complexity dynamics training an MLP on MNIST with weight decay. Robustness plots presented in Figure 28.